### v1

# Deformable Mamba for Wide Field of View Segmentation

Abstract
Wide-FoV cameras, like fisheye and panoramic setups, are essential for broader perception but introduce significant distortions in 180° and 360° images, complicating dense prediction tasks. For instance, existing MAMBA models lacking distortion-aware capacity cannot perform well in panoramic semantic segmentation. To address this problem, this work presents Deformable Mamba, a unified framework specifically designed to address imaging distortions within the context of panoramic and fisheye semantic segmentation. At the core is a decoder constructed with a series of Deformable Mamba Fusion (DMF) blocks, making the whole framework more deformable, efficient, and accurate, when handling extreme distortions. Extensive evaluations across five datasets demonstrate that our method consistently improves segmentation accuracy compared to the previous state-of-the-art methods tailored for specific FoVs. Notably, Deformable Mamba achieves a +2.5% performance improvement on the 360° Stanford2D3D dataset, and shows better results across FoVs from 60° to 360°.
Comparison of wide Field of View (wide FoV) segmentation.
Omni-directional performance across wide FoV from 60° to 360°.
Figure 1: Wide Field of View (wide-FoV) segmentation is investigated on 180° and 360° images. (a) Compared to other methods tailored for 90° or 180° data, our unified Deformable Mamba framework generalizes well on wide-FoV segmentation. (b) Besides, our methods yield consistently better results (mIoU %) across all different FoVs as compared to the baseline on the 360° Matterport3D dataset.
11footnotetext: Equal contribution.22footnotetext: Corresponding author (e-mail: jiaming.zhang@kit.edu).
1 Introduction
Dense image analysis like image segmentation [2, 58, 21] is a fundamental task in computer vision, forming a crucial component of numerous downstream vision-based applications [70]. At the same time, advancements in sensor technology have led to a variety of sensor types, each with unique characteristics that impact imaging. In recent years, diverse sensors such as pinhole cameras, fisheye cameras [64, 48] with 180° field of views (FoV), and even panoramic cameras [3, 1, 67] with 360° FoV have been increasingly adopted across vision tasks. However, differences in camera construction and sensor type result in distinct imaging characteristics, which can complicate traditional image processing [66]. Adapting algorithms to perform effectively with a specific FoV often requires considerable time and resources, involving method redesign, parameter tuning, and model retraining to address the unique challenges posed by each camera type.
Based on our experiments (Fig. Deformable Mamba for Wide Field of View Segmentation), we observe that straightforwardly applying or transferring existing specific models to new sensor types frequently leads to significant performance degradation. For instance, methods [21] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV. This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras, which complicates model generalization. Standard methods struggle to accurately capture the diverse object shapes and spatial distributions caused by such deformations. For instance, as shown in Fig. 2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits their adaptability. Previous methods have proposed solutions targeting specific sensor types, such as CNN-based and transformer-based semantic segmentation models [52, 31] optimized for fisheye images, as well as approaches [66, 67] tailored to mitigate distortion in panoramic segmentation. So far, to address diverse distortions across different camera types in a unified manner remains under-explored.
Figure 2: Comparison of camera imaging with pinhole (left) and fisheye (right) cameras with different wide-FoV. Pinhole cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
To address these challenges, we propose the Deformable Mamba model (DMamba for short), a versatile solution for image segmentation across varying wide-FoVs. In this work, a unified framework is created to effectively handling images from multiple camera types and sources, spanning FoVs from 180° to 360°, from indoor to outdoor scenarios, in synthetic and real-world imageries. At the core of our approach is a novel deformable decoder, which synergistically integrates State Space Models (SSMs) with deformable convolutions. Specifically, we leverage quadri-directional selective scan operations derived from VMamba [37] to capture global contextual information, while simultaneously employing deformable convolutions [77] to enhance feature representation. This dual-pathway desin effectively addresses the limitations of scanning mechanisms of Mamba [16] in visual applications while also endowing the model with the ability to perceive distortion. Our deformable decoder accommodates varied receptive fields, enabling seamless integration with CNN-based, transformer-based, and Mamba-based architectures. By leveraging this unified decoder, Deformable Mamba effectively captures features across disparate imaging perspectives, making it an adaptable solution for segmentation tasks across a broad spectrum of camera types.
To evaluate the effectiveness of Deformable Mamba, we conducted extensive experiments across five datasets for semantic segmentation, including three 360° (SynPASS [67], Stanford2D3D [1], Matterport3D [3]) and two 180° (WoodScape [64], SynWoodScape [48]) datasets. The data encompass a variety of scenarios including 180° and 360° images, indoor and outdoor environments, and both synthetic and real-world data. Our unified framework performs consistently well across various datasets. Notably, on 360° Stanford2D3D dataset, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over the pinhole Mamba model. In 180° SynWoodScape dataset, it similarly surpassed prior best-performing models, achieving an increase of ${+}2.2\%$ in mean Intersection over Union (mIoU). Further ablation studies, including comprehensive omni-directional comparisons, consistently demonstrated the model’s effectiveness in handling wide-FoVs, showing its potential as a general solution for dense image analysis across diverse sensor types. We hope this work opens up new possibilities for unifying wide-FoV segmentation.
To summarize, our contributions are as follows.
  • We propose a Mamba-based framework termed Deformable Mamba, with the aim at unifying distortion-aware wide-FoV segmentation.
  • A deformable decoder constructed with Deformable Mamba Fusion (DMF) module is specifically designed to boost the performance of mamba-based models in the wide-FoV segmentation task.
  • Extensive experiments on five datasets across a variety of indoor and outdoor, synthetic and real-world wide-FoV scenes, demonstrate the state-of-the-art performance of our method in wide-FoV segmentation.
2 Related Work
2.1 Wide Field of View Segmentation
Wide Field-of-View (Wide-FoV) semantic segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics. However, directly applying pinhole models [39, 2, 5, 15, 26, 65, 65, 36, 34, 72, 50, 58, 9, 19, 6, 70] to wide-FoV images often results in downgraded performance due to object deformations and image distortions. Previous work [61, 62] propose the Panoramic Annular Semantic Segmentation (PASS) framework using a panoramic annular lens system.  [27] introduces panoramic panoptic segmentation for panoramic views. In [66, 67], deformable modules are proposed to enhance the distortion perception capability in transformers. Zheng et al. [71] introduce the Open Panoramic Segmentation (OPS) task in an open-vocabulary context. Apart from panoramas, previous methods [31, 11] revolve around multi-task on fisheye images, while a pretraining paradigm [42] and semi-supervised approaches [45, 43] are introduced to address the challenges of fisheye data. Manzoor et al. [40] substitutes the conventional convolutional modules in U-Net [47] and residual U-Net [46] with restricted deformable convolution (RDC) [11], imparting fisheye-distortion awareness to the model. Departing from these approaches, we are the first to propose a unified model applicable to all wide-FoV inputs, including panoramic and fisheye images.
2.2 State Space Model
CNNs [30] have achieved remarkable success in computer vision through effective local feature extraction and hierarchical representation learning capabilities, while Transformers [54] excel at modeling long-range dependencies despite their quadratic computational complexity. Recently, the state-space-model (SSM) based Mamba [16] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, ushering in a new paradigm for long sequence modeling. The emergence of Mamba in sequence modeling has naturally extended to vision tasks, Zhu et al. [75] adopt the patch embedding from Vision Transformer (ViT) [12] to project 2D image patches into sequential tokens compatible with Mamba architecture. To address the inherent discrepancy between image and sequence representations, a quadri-directional scanning and merging mechanism termed SS2D is introduced in [37] to supersede Mamba’s canonical 1D scan operation. Subsequent work [25, 44, 24, 32, 69, 63, 4] introduce various scanning methods to explore variations of Mamba in visual representations. However, the study of using SSM for wide-FoV imagery is under-explored. Therefore, for the first time, we propose deformable Mamba to address both panoramic and fisheye semantic segmentation.
3 Methodology
3.1 Preliminary
Image Distortion. As shown in the left side of Fig. 2, the point $A$ in the world coordinates undergoes extrinsic transformation to camera coordinates by Eq. (1), where $R$ and $T$ denote the rotation matrix and translation vector, respectively. Following Eq. (2), points in the camera coordinates are projected onto the image plane, with $f_{x}$ and $f_{y}$ representing the focal lengths and $c_{x},c_{y}$ defining the principal point. While straight lines are preserved in pinhole-approximate narrow-FoV cameras, only points within the sensor’s FoV (red box in Fig. 2) are imaged, excluding points like $B^{\prime}$ that fall outside this region. Narrow-FoV cameras follow perspective projection principles where object size in the image varies inversely with distance between object and camera, introducing moderate distortion.
$\begin{bmatrix}\mathbf{X}_{c}\\ \mathbf{Y}_{c}\\ \mathbf{Z}_{c}\end{bmatrix}=\mathbf{R}\begin{bmatrix}\mathbf{X}_{w}\\ \mathbf{Y}_{w}\\ \mathbf{Z}_{w}\end{bmatrix}+\mathbf{T},$ (1)
$\begin{bmatrix}\mathbf{x^{\prime}}\\ \mathbf{y^{\prime}}\\ 1\end{bmatrix}=\begin{bmatrix}\mathbf{f}_{x}&0&\mathbf{c}_{x}\\ 0&\mathbf{f}_{y}&\mathbf{c}_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\frac{\mathbf{X}_{c}}{\mathbf{Z}_{c}}\\ \frac{\mathbf{Y}_{c}}{\mathbf{Z}_{c}}\\ 1\end{bmatrix}.$ (2)
In contrast, wide-FoV cameras like fisheye cameras employ complex projection models involving light refraction. Kannala et al. [28] propose a generalized projection model for wide-FoV cameras, the distance between the image point and the principal point $r$ is approximated by:
$\mathbf{r}(\mathbf{\theta})=\mathbf{k_{1}}\mathbf{\theta}+\mathbf{k_{2}}\mathbf{\theta}^{3}+\mathbf{k_{3}}\mathbf{\theta}^{5}+\mathbf{k_{4}}\mathbf{\theta}^{7}+\mathbf{k_{5}}\mathbf{\theta}^{9},$ (3)
where $\theta$ is the angle between the principal axis and the incoming ray, $k_{i},i\in\{1,2,3,4,5\}$ are the parameters of wide-FoV camera model. While these cameras capture wider FoV, they suffer from more severe geometric distortions compared to the perspective projection of narrow-FoV cameras. As shown in the right side of Fig. 2, barrel distortion in wide-FoV cameras causes straight lines to appear curved outward, particularly pronounced near image edges while the center remains relatively undistorted. This distortion manifests as radially-increasing displacement from the center, resulting in progressive spatial compression toward the image periphery.
State Space Model. State space model (SSM), which describes the dynamics of a system, can be formulated as:
$\begin{split}\mathbf{h}^{\prime}(\mathbf{t})&=\mathbf{A}\mathbf{h}(\mathbf{t})+\mathbf{B}\mathbf{x}(\mathbf{t}),\\ \mathbf{y}(\mathbf{t})&=\mathbf{C}\mathbf{h}(\mathbf{t}),\\ \end{split}$ (4)
where $x(t)$ and $y(t)$ denote the inputs and outputs of the system, respectively, $h(t)$ is the inner states of the system, $A,B,C$ are the system parameters. Gu et al.  [18] propose Structured State Space Sequence Model (S4) by initializing the parameters $A$ using Hippo Matrix (5) and discretizing parameters A and B by step size $\Delta$.
$\mathbf{A}_{nk}=-\begin{cases}(2\mathbf{n}+1)^{1/2}(2\mathbf{k}+1)^{1/2}&\text{if }\mathbf{n}>\mathbf{k}\\ \mathbf{n}+1,&\text{if }\mathbf{n}=\mathbf{k}\\ 0,&\text{if }\mathbf{n}<\mathbf{k}.\end{cases}$ (5)
In Mamba [16], a selective scan mechanism is proposed to further extend the SSM from time-invariant to time-variant, while also enabling parallel training. Observations. Based on HiPPO theory [17], Mamba [16] is inherently a sequence model, where the output of the current token only depends on preceding tokens, without considering subsequent ones - a limitation that becomes problematic for 2D images with inherent spatial structures. Unlike Transformer’s [54] mechanism in ViT [12], where attention computation for each patch considers all patches globally, Mamba processes image patches sequentially.
To address this limitation, previous work [37, 75, 4, 76, 63] have explored different scanning patterns, including bidirectional or quadri-directional scanning approaches, partially addressing the challenges posed by Mamba’s unidirectional scanning nature in visual tasks. However, these multi-directional scanning approaches face another challenge. As illustrated in Fig 3., consider an input image embedded into 9 patches (labeled 1-9). When modeling patch 5 using SS2D’s [37] quadri-directional scanning mechanism, the scanning sequences are:
  • (1) Left-Right: 1 - 2 - 3 - 4 - 5
  • (2) Right-Left: 9 - 8 - 7 - 6 - 5
  • (3) Top-Down: 1 - 4 - 7 - 2 -5
  • (4) Down-Top: 9 - 6 - 3 - 8 - 5
Figure 3: Visualization of SS2D quadri-directional scanning.
Patch 1 and patch 3 maintain equal distances from patch 5, demonstrating uniform spatial distribution. As illustrated in sequences (1) (3) (4), the scanning trajectory from patch 1 to patch 5 traverses double the path length compared to the direct route between patch 3 and patch 5, despite their equivalent spatial distances. Since the Mamba model learns relationships between serialized tokens, the contribution of earlier patches in the scanning path to the current patch decreases as their distance from the current patch increases. Thus, along this scanning path, we infer that patch 1 is more important to patch 5 than patch 3. This priority is introduced by the scanning mechanism, but it is not a reasonable assumption. Additionally, when computing for other target patches, this priority changes based on the target’s position.
Figure 4: Overview of the Deformable Mamba (DMamba) framework. Given wide-FoV images (180° or 360°), the features extracted by an encoder, are fused by the proposed Deformable Mamba Decoder, which constructed by four Deformable Mamba Fusion (DMF) modules.
3.2 Overall Architecture
An overview of our Deformable Mamba architecture is presented in Fig. 4, which illustrates how we make Mamba-based models deformable and available for wide-FoV semantic segmentation. It mainly follows the Encoder-Decoder architecture. The backbone selection is flexible for different methods like CNN-, transformer- or Mamba-based architectures. In this work, we explore with VMamba [37] as the backbone and our proposed Deformable Mamba Decoder. In our framework, there are two model sizes: DMamba-M (Mini) and DMamba-T (Tiny).
More specifically, given the wide-FoV input image in the shape of $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$. The features extracted by the encoder undergo a hierarchical fusion process in the decoder.
3.3 Deformable Mamba Decoder
The Deformable Mamba Decoder comprises four fusion and upsampling modules, where we introduce a novel Deformable Mamba Fusion (DMF) module for multi-scale feature integration. The initial DMF block processes dual inputs of ${\frac{H}{32}\times\frac{H}{32}\times 768}$ feature maps derived from the final encoder layer. The fused features are then processed through an upsampling module that simultaneously reduces channel dimensionality and increases spatial resolution, followed by channel mixing via an MLP layer. Following [49, 68, 35, 60], we utilize Pixel Shuffle [49] as the upsampling operation. In the subsequent stages, the second through fourth DMF blocks implement skip connections, facilitating interaction between the previously fused and upsampled features and their corresponding encoder features at matching scales. This progressive fusion strategy enables comprehensive multi-scale feature integration. The decoder’s output undergoes final processing through a fully connected layer and class mapping to generate the segmentation predictions.
3.4 Deformable Mamba Fusion
To address the aforementioned limitations of Mamba’s scanning mechanism in visual tasks (Sec. 3.1), we propose the DMF module, which simultaneously adapts the features extracted by the Mamba-based encoder and integrates multi-scale features. As illustrated in the lower part of Fig. 4, the DMF module takes two inputs: one is the feature $\mathbf{D}_{j}$ with dimensions $c\times h\times w$ output from the upper decoder layer, and the other is the corresponding encoder features $\mathbf{E}_{i}$ at matching scales, where $i$ and $j$ satisfy the relationship $i+j=5$. We employ the Cross-Scan [37] that scans $\mathbf{D}_{j}$ in four directions: (1) Left to Right, (2) Right to Left, (3) Top to Down, and (4) Down to Top, utilizing Mamba model’s global perception to generate four feature maps with resolution $c\times h\times w$ Following Cross-Merge [37], we first reverse features (2) and (4) along their last dimension and add them to (1) and (3) respectively, then transpose the $h$ and $w$ dimensions of feature (3) before adding it to (1).
Inspired by previous works [10, 77, 56, 59], we propose a simple yet effective approach, utilizing deformable convolution [77] to enhance the features extracted by the Mamba model. Specifically, for a feature point in the encoder feature map $\mathbf{E}_{i}$, we attend to $\mathbf{K}$ neighboring feature points, allowing the model to learn offsets that compensate for the biases introduced by the scanning process. Additionally, convolution operations on these selected regions provide a more localized attention mechanism, further enhancing the visual representation capability of Mamba-based models in visual tasks. The mathematical formulation is as follows:
$\mathbf{y}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{x}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (6)
where $\mathbf{p}$ denotes the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively. And $\mathbf{m_{k}}$ is a learnable modulation scalar, which helps to modulate the input feature amplitudes from different spatial locations. For example, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
The outputs from these two branches are concatenated and fed into a convolutional layer for feature fusion. This simple yet effective architecture successfully adapts the features extracted by the Mamba model, efficiently integrates multi-scale features, while endowing the entire decoder with deformable capabilities. Experiments on multiple datasets demonstrate both the effectiveness and efficiency of our proposed model.
4 Experiment Results
4.1 Datasets
To study the unification of wide-FoV segmentation, we conduct experiments on three 360° and two 180° datasets.
Stanford2D3D [1] consists of 1,413 indoor 360° images with a resolution of $512{\times}1024$, encompassing 13 distinct object categories.
Matterport3D [3] has 10,615 indoor 360° images. Following [53, 22], 7,829 of panoramas with a resolution of 1024×2048 are used for training and 772 for evaluation.
SynPASS [67] is a virtual outdoor 360° dataset from CARLA [13]. It contains 9,080 images with a resolution of $1024{\times}2048$, covering various weather conditions (cloudy, foggy, rainy, sunny) and illumination (daytime, nighttime).
WoodScape [64] consists of 10K 180° images captured by four surround view cameras. The semantic labels have 9 categories. Since the test split is not available, 8,234 released images are reallocated into training and evaluation sets with an 80% to 20% split.
SynWoodScape [48] is a synthetic version of the surround-view dataset. The released v0.1.1 version contains 2,000 180° surround-view fisheye RGB images. We split them into training and validation sets with an 80% to 20 % ratio.
4.2 Implementation Details
In this work, our models are trained and validated on three 360° datasets and two 180° datasets, covering both indoor and outdoor scenarios, as well as real and synthetic environments, using 4 A100 GPUs. For all experiments, we employed cross-entropy as the loss function. To validate the effectiveness of our framework, we deliberately avoided using any auxiliary losses. The models were optimized using AdamW with an initial learning rate of $6e^{-5}$ and weight decay of 0.01. For Stanford2D3D and SynWoodscape, we trained for 80K iterations with a batch size of 2 per GPU. For Matterport3D, SynPASS, and Woodscape, we extended the training to 160K iterations, maintaining the same batch size of 2 per GPU. The learning rate schedule consisted of a linear warm-up for the first 1.5K iterations, followed by a polynomial decay with a power of 0.9.
4.3 Quantitative Results
Given that previous works were largely constrained to specific Field-of-View (FoV) settings, which poses challenges for fair cross-dataset evaluation. Therefore, we selected representative state-of-the-art models as our baselines, including SegNeXt [21], which achieves superior performance on pinhole images, and 360SFUDA [73], which is specifically designed for panoramic scenarios. To ensure fair comparison, we carefully match the parameter counts across all implementations and conduct extensive evaluations on the aforementioned datasets with varying FoV characteristics.
Table 1: Results on 360° segmentation datasets of Stanford2D3D (S2D3D) fold-1 and 360FV-Matterport (MP3D). denotes our re-implementation.
  Method Backbone mIoU
S2D3D
MP3D
Tangent [14] (CVPR20) ResNet-101 45.6 -
HoHoNet [51] (CVPR21) ResNet-101 52.0 44.1
CBFC [74] (CVPR23) ResNet-101 52.2 -
Trans4PASS [66] (CVPR22) Trans4PASS 52.1 41.9
Trans4PASS+ [67] (PAMI24) Trans4PASS+ 54.0 42.6
SegFormer-B2 [58] (NeurIPS21) MiT-B2 51.9 45.5
360BEV [53] (WACV24) MiT-B2 54.3 46.3
SGAT4PASS [33] (IJCAI23) MiT-B2 56.4 -
360SFUDA [73] (CVPR24) MixT-B2 54.9 46.7
360SFUDA [73] (CVPR24) MixT-B3 56.9 47.9
SegNeXt [21] (NeurIPS22) MSCAN-B 57.0 48.4
VMamba [37] (NeurIPS24) VMamba-T 56.8 47.6
DMamba-M (Ours) VMamba-T 59.3 48.7
DMamba-T (Ours) VMamba-S 60.2 49.3
Results on 360° Stanford2D3D. Table 1 shows superior performance of our method compared to both panoramic [14, 51, 74, 66, 67, 53, 33, 73] and pinhole-based [21, 37] methods. Specifically, our model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines.
Notably, compared to the VMamba-T + UperNet [57], our lightweighted DMamba-M with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing parameter count by 30% and FLOPs by 80%. Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba-T + UperNet, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
Table 2: Results on 360° segmentation dataset SynPASS. denotes our re-implementation.
  Method
mIoU
Building
Fense
Other
Pedestrian
Pole
RoadLine
Road
SideWalk
Vegetation
Vehicles
Wall
TrafficSign
Sky
Ground
Bridge
RailTrack
GroundRail
TrafficLight
Static
Dynamic
Water
Terrain
PVT-T [55] (CVM22) 32.37 74.83 19.94 00.24 21.82 13.15 62.59 93.14 49.09 67.27 46.44 01.69 09.63 96.09 00.18 02.64 08.81 61.11 14.09 12.04 00.99 05.05 51.33
PVT-S [55] (CVM22) 32.68 78.02 27.12 00.27 23.48 16.51 59.81 92.87 50.21 66.22 43.50 01.12 08.67 96.34 00.44 00.15 02.82 63.88 13.78 15.15 01.58 08.78 48.29
SegFormer-B1 [58] (NeurIPS21) 37.36 78.24 20.59 00.00 38.28 21.09 68.72 94.50 59.72 68.43 67.51 00.83 09.86 96.08 00.56 01.38 20.79 69.59 23.38 19.91 01.38 08.97 52.07
SegFormer-B2 [58] (NeurIPS21) 37.24 79.25 23.58 00.00 40.01 20.14 65.28 92.80 46.92 68.64 77.45 01.42 15.00 96.33 00.57 00.58 02.68 67.60 25.89 20.80 01.99 20.92 51.53
Trans4PASS-T [66] (CVPR22) 38.53 79.17 28.18 00.13 36.04 23.69 69.16 95.51 61.71 69.77 71.12 01.53 16.98 96.50 00.56 01.60 15.22 70.48 26.03 23.11 02.08 09.24 49.77
Trans4PASS-S [66] (CVPR22) 38.57 80.02 24.56 00.07 41.49 25.23 72.00 95.89 59.88 69.07 77.08 01.04 13.72 96.69 00.67 00.73 05.60 72.56 25.93 22.45 02.78 08.34 52.65
Trans4PASS+(T) [67] (PAMI24) 39.42 79.63 24.45 00.21 44.23 26.71 70.32 95.86 61.80 69.25 78.85 01.09 13.81 97.12 00.91 03.48 19.32 72.44 21.08 25.56 02.67 05.03 53.20
Trans4PASS+(S) [67] (PAMI24) 40.72 80.91 20.78 00.23 45.36 24.08 72.51 96.79 67.15 70.46 81.39 04.28 26.19 97.21 01.24 01.74 16.56 67.08 28.64 23.68 03.35 08.48 57.57
SegNeXt-B [21] (NeurIPS22) 39.90 82.62 16.06 00.12 43.21 26.15 71.69 97.10 60.64 67.56 82.56 00.32 24.43 97.23 02.03 00.00 43.23 47.90 31.61 26.50 03.63 10.56 42.68
360SFUDA-B2 [73] (CVPR24) 40.37 82.69 22.27 00.16 41.86 28.19 74.98 97.13 62.22 70.05 81.55 04.21 21.09 97.45 01.06 00.49 31.77 51.08 32.15 24.88 03.59 10.02 49.23
360SFUDA-B3 [73] (CVPR24) 41.33 82.91 32.33 00.23 41.87 27.21 75.13 97.37 62.90 69.34 82.95 01.66 26.03 97.56 00.95 00.10 41.39 64.02 32.30 26.74 03.48 16.76 48.14
DMamba-M(Ours) 41.72 83.19 21.52 00.21 48.15 32.20 75.8 97.51 61.00 67.74 81.86 01.06 31.30 94.47 01.37 00.68 29.63 49.41 36.35 30.99 04.08 20.54 45.27
DMamba-T(Ours) 42.78 84.24 22.57 00.25 49.20 33.25 76.85 98.56 62.05 68.79 82.91 02.11 32.35 98.52 02.42 01.73 30.68 50.46 37.40 32.04 05.85 21.59 46.32
Results on 360° Matterport3D.
While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of 0.3 mIoU over the previous SOTA model SegNeXt as shown in Table 1. Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
Results on 360° SynPASS. Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the SynPASS dataset. It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset. This discrepancy underscores the sensitivity of certain architectures to distribution shifts. In contrast, our method, as shown in Table 2, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods. Specifically, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic. For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
Table 3: Results on 180° segmentation dataset WoodScape. denotes our re-implementation.
  Method Mutli-task Supervised mIoU
MeanTeacher [52] (NeurIPS17) 54.20
CPS [7] (CVPR21) 60.31
CPS with CutMix [7] (CVPR21) 62.47
FishSegSSL [43] (J. Imaging24) 64.81
OmniDet [31] (RA-L21) 72.50
    + DTP [20] (ECCV18) 75.80
    + GradNorm [8] (PMLR18) 75.90
    + Uncertainity [29] (CVPR18) 76.10
    + VarNorm [31] (RA-L21) 76.60
SegNeXt-B [21] (NeurIPS22) 82.32
360SFUDA-B2 [73] (CVPR24) 81.74
360SFUDA-B3 [73] (CVPR24) 82.43
DMamba-M (Ours) 83.21
DMamba-T (Ours) 83.90
Results on 180° WoodScape. In Table 3, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [64]. Semi-supervised methods such as MeanTeacher [52], CPS [7], and FishSegSSL [43] show lower performance due to the limited data. Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [31] model.
Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively. The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
Table 4: Results on 180° segmentation dataset SynWoodScape. denotes our re-implementation.
             Method            mIoU
           SegNeXt-B [21] (NeurIPS22)            64.86
           360SFUDA-B2 [73] (CVPR24)            66.56
           360SFUDA-B3 [73] (CVPR24)            67.47
           DMamba-M (Ours)            69.20
           DMamba-T (Ours)            69.62
Results on 180° SynWoodScape. Due to the lack of baseline comparisons on SynWoodScape, we limited our evaluation to SegNeXt and 360SFUDA. Similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic 180° data. This further confirms the challenge of achieving consistently high performance across diverse datasets. In comparison, our DMamba-M of similar scale outperforms them by ${+}4.36\%$ and ${+}1.73\%$ mIoU, respectively, achieving an improved segmentation accuracy of 69.20 mIoU on SynWoodScape.
These consistent improvements across all five wide-FoV datasets demonstrate the effectiveness of our proposed framework in unifying wide-FoV segmentation.
Figure 5: Visualization of the wide-FoV segmentation results. From left to right: one 360° and two 180° segmentation.
4.4 Qualitative Results
To demonstrate the improved performance of our model over Mamba-based models, we compared the visualization results of DMamba and VMamba on the Stanford2D3D and SynWoodScape. As discussed in Sec. 3.1 and  3.4, the scanning approach in Mamba-based models introduces biases in visual tasks that limit their effectiveness in dense prediction tasks. These biases increase randomness and reduce the model’s ability to capture local spatial relationships, which is particularly evident in the recognition and segmentation of smaller objects. As shown in Fig. 5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings. In contrast, our proposed DMamba effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
Table 5: Ablation study for analysis of backbones (BB) and decoders (Dec) on the 360° dataset Stanford2D3D.
[…]
BB
Dec
Total
BB
Dec
Total
mIoU
mAcc
ResNet-UperNet ResNet50 [23] UperHead [57] 44.3 23.5 40.5 64.0 22.9 250.7 273.6 53.34 62.42
ResNet-UperFormer ResNet50 [23] MusterHead [60] 14.3 23.5 203.1 226.6 22.9 211.5 234.4 54.39 63.27
ResNet-CGRSeg ResNet50 [23] CGRHead [41] 58.0 23.5 282.6 306.1 22.9 31.4 54.3 54.12 62.94
ResNet-DMamba (Ours) ResNet50 [23] DMambaHead 27.9 23.5 77.3 100.8 22.9 38.8 61.7 57.22 66.32
Swin-UperNet Swin-T [38] UperHead [57] 34.9 27.5 31.5 59.0 25.6 211.5 237.1 53.82 63.06
Swin-UperFormer Swin-T [38] MusterHead [60] 30.2 27.5 19.1 46.6 25.6 21.9 47.5 52.36 54.54
Swin-CGRSeg Swin-T [38] CGRHead [41] 52.2 27.5 40.6 68.1 25.6 5.5 31.1 54.26 62.90
Swin-DMamba (Ours) Swin-T [38] DMambaHead 37.3 27.5 11.2 38.7 25.6 6.5 32.1 55.09 65.30
VMamba-UperNet VMamba-T [37] UperHead [57] 35.0 29.5 31.5 61.0 25.3 206.9 232.2 56.80 65.01
VMamba-UperFormer VMamba-T [37] MusterHead [60] 29.8 29.5 19.1 48.6 25.3 21.4 46.7 58.12 66.97
VMamba-CGRSeg VMamba-T [37] CGRHead [41] 42.2 29.5 40.5 70.0 25.3 5.0 30.3 56.45 65.07
DMamba-M (Ours) VMamba-T [37] DMambaHead 40.0 29.5 11.1 40.6 25.3 6.0 31.3 59.30 68.99
DMamba-T (Ours) VMamba-S [37] DMambaHead 17.9 49.4 11.1 60.5 44.7 6.0 50.7 60.18 71.68
4.5 Ablation Study
Effect of Backbone. In Table 5, we ablate three different kinds of commonly-used model backbones, namely CNN-based ResNet [23], ViT-based Swin Transformer [38] and Mamba-based VMamba [37]. For the sake of a fair comparison, we choose the backbones with a comparable number of learnable parameters and FLOPs. According to Table 5 it can be observed that the models equipped with VMamba [37] outperform other models consistently while CNN-based ResNet50 and ViT-based Swin Transformer achieve similar performance. This conclusion can be validated by employing the same decoder, such as UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead, along with a consistent model backbone to perform the segmentation task. Notably, the combination of the VMamba [37] backbone with DMambaHead, resulting in our Deformable Mamba model, demonstrates superior performance compared to other variant combinations by a maximum of over $+6.9\%$ mIoU.
Effect of Decoder. Apart from the backbones, we also ablate the decoders. For every backbone, we conduct four experiments with different decoders, i.e., UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead. All results are reported in Table 5. Using the same model backbone, our proposed DMambaHead outperforms other widely-used decoders by a $+3\%$ mIoU while a complexity degradation of $+20\%$, making the Deformable Mamba and its variants extremely practical in real-time wide-FoV accurate segmentation. It’s worth noting that the DMambaHead with VMamba-T [37] backbone achieves a drop of FLOPs from $232.2$G to $31.3$G, a drop of the learnable parameters from $61$M to $40.6$M while the performance of Deformable Mamba increases from a $+56.8\%$ to $+59.3\%$ mIoU compared with the VMamba-T [37] and UperHead [57] combination. The reason behind this is that our proposed DMambaHead is specifically designed for wide-FoV segmentation. Our deformable design enables the model to focus on the distortion introduced by the wide-FoV fisheye or panoramic images. This ablation proves that not only the VMamba [37] backbone but also the proposed DMambaHead contributes to the superior performance and efficiency of the Deformable Mamba model.
(a) Results on Matterport3D.
(b) Results on Stanford2D3D.
Figure 6: Analysis of segmentation with different wide FoVs.
Analysis of FoVs. To further analyze the wide-FoV segmentation, we conduct omni-directional analysis. As shown in Fig. 6, there are six different FoVs, ranging from 60° to 360°. Compared to the baseline VMamba [37], our DMamba methods obtain relative better results on all different FoVs. Note that the 60° performance of all methods are lower than other degrees on the Stanford2D3D dataset. One reason is that the object distribution and room layout in the 60° direction are different between Stanford2D3D and the Matterport3D datasets. Nonetheless, our Mini- and Tiny-version methods show consistent improvements over the baseline on both datasets in different FoVs.
5 Conclusion
In this work, we are the first to explore the unification of the wide field of view (FoV) segmentation tasks, including 180° and 360° image segmentation. Most existing models are tailored for pinhole camera images, making it challenging to extend their success to wide-FoV domains due to the severe image distortion and object deformation. Unlike previous approaches that require a specific architecture for each FoV or dataset, we propose a unified framework, Deformable Mamba, built on the emerging State Space Model (SSM) mechanism. Our framework is constructed with a flexible deformable decoder, which can be easily adapted to various models and datasets. Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our unified framework effectively addresses the unique challenges of wide-FoV segmentation.
Limitations and Future Work. Using large language models (LLMs) for unifying the wide-FoV segmentation is a potential solution but not discussed in this work, while LLMs can further include the segmentation methods that use vision-language architectures. Besides, cross-domain evaluation, e.g., transferring from a synthetic to a real wide-FoV scene, or from a narrow-FoV camera to a wide-FoV camera, will be further explored in our future work.

### v2

@@
 Abstract
-Wide-FoV cameras, like fisheye and panoramic setups, are essential for broader perception but introduce significant distortions in 180° and 360° images, complicating dense prediction tasks.
-For instance, existing MAMBA models lacking distortion-aware capacity cannot perform well in panoramic semantic segmentation.
-To address this problem, this work presents Deformable Mamba, a unified framework specifically designed to address imaging distortions within the context of panoramic and fisheye semantic segmentation.
-At the core is a decoder constructed with a series of Deformable Mamba Fusion (DMF) blocks, making the whole framework more deformable, efficient, and accurate, when handling extreme distortions.
-Extensive evaluations across five datasets demonstrate that our method consistently improves segmentation accuracy compared to the previous state-of-the-art methods tailored for specific FoVs.
-Notably, Deformable Mamba achieves a +2.5% performance improvement on the 360° Stanford2D3D dataset, and shows better results across FoVs from 60° to 360°.
-Comparison of wide Field of View (wide FoV) segmentation.
-Omni-directional performance across wide FoV from 60° to 360°.
-Figure 1: Wide Field of View (wide-FoV) segmentation is investigated on 180° and 360° images.
-(a) Compared to other methods tailored for 90° or 180° data, our unified Deformable Mamba framework generalizes well on wide-FoV segmentation.
-(b) Besides, our methods yield consistently better results (mIoU %) across all different FoVs as compared to the baseline on the 360° Matterport3D dataset.
+Recent advancements in the Mamba architecture, with its linear computational complexity, being a promising alternative to transformer architectures suffering from quadratic complexity.
+While existing works primarily focus on adapting Mamba as vision encoders, the critical role of task-specific Mamba decoders remains under-explored, particularly for distortion-prone dense prediction tasks.
+This paper addresses two interconnected challenges: (1) The design of a Mamba-based decoder that seamlessly adapts to various architectures (e.g., CNN-, Transformer-, and Mamba-based backbones), and (2) The performance degradation in decoders lacking distortion-aware capability when processing wide-FoV images (e.g., 180° fisheye and 360° panoramic settings).
+We propose the Deformable Mamba Decoder, an efficient distortion-aware decoder that integrates Mamba’s computational efficiency with adaptive distortion awareness.
+Comprehensive experiments on five wide-FoV segmentation benchmarks validate its effectiveness.
+Notably, our decoder achieves a +2.5% performance improvement on the 360° Stanford2D3D segmentation benchmark while reducing 72% parameters and 97% FLOPs, as compared to the widely-used decoder heads.
 11footnotetext: Equal contribution.22footnotetext: Corresponding author (e-mail: jiaming.zhang@kit.edu).
+(a) Comparison across narrow- and wide-FoV semantic segmentation.
+(b) Performance and efficiency comparison of different decoders on Stanford2D3D [1] panoramic dataset with the same backbone [34].
+Figure 1: Our deformable Mamba (a) achieves better results across wide-FoV semantic segmentation while (b) maintaining parameter and computational efficiency.
 1 Introduction
@@
 In recent years, diverse sensors such as pinhole cameras, fisheye cameras [62, 45] with 180° field of views (FoV), and even panoramic cameras [3, 1, 65] with 360° FoV have been increasingly adopted across vision tasks.
-However, differences in camera construction and sensor type result in distinct imaging characteristics, which can complicate traditional image processing [66].
-Adapting algorithms to perform effectively with a specific FoV often requires considerable time and resources, involving method redesign, parameter tuning, and model retraining to address the unique challenges posed by each camera type.
 Based on our experiments (Fig.
-Deformable Mamba for Wide Field of View Segmentation), we observe that straightforwardly applying or transferring existing specific models to new sensor types frequently leads to significant performance degradation.
-For instance, methods [21] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV.
-This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras, which complicates model generalization.
-Standard methods struggle to accurately capture the diverse object shapes and spatial distributions caused by such deformations.
-For instance, as shown in Fig.
-2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits their adaptability.
-Previous methods have proposed solutions targeting specific sensor types, such as CNN-based and transformer-based semantic segmentation models [52, 31] optimized for fisheye images, as well as approaches [66, 67] tailored to mitigate distortion in panoramic segmentation.
-So far, to address diverse distortions across different camera types in a unified manner remains under-explored.
-Figure 2: Comparison of camera imaging with pinhole (left) and fisheye (right) cameras with different wide-FoV.
-Pinhole cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
-To address these challenges, we propose the Deformable Mamba model (DMamba for short), a versatile solution for image segmentation across varying wide-FoVs.
+1(a)), we observe that directly applying models designed for narrow-FoV data to wide-FoV data often leads to a suboptimal solution.
+For instance, SegNext [20] specifically designed for narrow-FoV pinhole cameras typically cannot generalize well when used for analyzing fisheye images with 180° FoV or panoramic images with 360° FoV.
+This is primarily due to the high degree of image distortion and object deformation present in wide-FoV cameras.
+As shown in Fig.
+2, the altered pixel arrangement in wide-FoV images might lead to misclassifications and limits adaptability of narrow-FoV models [2, 5, 7, 69, 47].
+To address this, previous methods have spent a large effort in architecture design to fit specific sensor types, such as CNN-based and transformer-based models [49, 29] for fisheye images, as well as approaches [64, 65] tailored to mitigate distortion in panoramic segmentation.
+However, most existing methods deeply couple distortion-aware capability with the entire model (both encoder and decoder).
+The drawbacks are two-fold: (1) These models are often designed for specific FoVs, yet hard to adapt to varying distortion characteristics.
+For example, narrow-FoV model SegNext [20] and wide-FoV model 360SFUDA [70] cannot consistently perform well across different wide-FoV segmentation (Fig.
+1(a)).
+(2) The tight coupling with the overall architecture limits the reusability of large-scale pretrained backbones, as any modification requires retraining.
+Therefore, in this work, we focus on flexible decoder designs and constrain distortion awareness to the decoder level, yielding greater flexibility in utilizing diverse pretrained backbones, striking a tradeoff between adaptability and modularity.
+Figure 2: Comparison of camera imaging with narrow-FoV (left) and wide-FoV (right) cameras.
+Narrow-FoV cameras maintain geometric fidelity but limited coverage, whereas wide-FoV cameras offer expansive scene capture while introducing substantial geometric distortions.
+Furthermore, we mainly utilize the Mamba [17] mechanism due to its benefit of linear computational complexity.
+However, most previous works [72, 34, 58, 24, 30, 4, 73] have focused on Mamba-based encoders, research on corresponding decoders for downstream tasks remains limited.
+We contend that decoders are equally crucial as encoders in the context of dense prediction tasks.
+As illustrated in Fig.
+1(b), a widely used decoder UperHead [54] excels at multi-scale feature fusion but suffers from excessive computational overhead.
+Alternative methods [57, 38] applied lightweight CNN- and transformer-based decoders to reduce computational cost.
+However, they suffer notable performance degradation when processing wide-FoV data.
+To address these challenges, we propose the Deformable Mamba Decoder , an efficient distortion-aware decoder for image segmentation across varying wide-FoVs.
 In this work, a unified framework is created to effectively handling images from multiple camera types and sources, spanning FoVs from 180° to 360°, from indoor to outdoor scenarios, in synthetic and real-world imageries.
-At the core of our approach is a novel deformable decoder, which synergistically integrates State Space Models (SSMs) with deformable convolutions.
-Specifically, we leverage quadri-directional selective scan operations derived from VMamba [37] to capture global contextual information, while simultaneously employing deformable convolutions [77] to enhance feature representation.
-This dual-pathway desin effectively addresses the limitations of scanning mechanisms of Mamba [16] in visual applications while also endowing the model with the ability to perceive distortion.
-Our deformable decoder accommodates varied receptive fields, enabling seamless integration with CNN-based, transformer-based, and Mamba-based architectures.
-By leveraging this unified decoder, Deformable Mamba effectively captures features across disparate imaging perspectives, making it an adaptable solution for segmentation tasks across a broad spectrum of camera types.
-To evaluate the effectiveness of Deformable Mamba, we conducted extensive experiments across five datasets for semantic segmentation, including three 360° (SynPASS [67], Stanford2D3D [1], Matterport3D [3]) and two 180° (WoodScape [64], SynWoodScape [48]) datasets.
-The data encompass a variety of scenarios including 180° and 360° images, indoor and outdoor environments, and both synthetic and real-world data.
-Our unified framework performs consistently well across various datasets.
-Notably, on 360° Stanford2D3D dataset, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over the pinhole Mamba model.
-In 180° SynWoodScape dataset, it similarly surpassed prior best-performing models, achieving an increase of ${+}2.2\%$ in mean Intersection over Union (mIoU).
-Further ablation studies, including comprehensive omni-directional comparisons, consistently demonstrated the model’s effectiveness in handling wide-FoVs, showing its potential as a general solution for dense image analysis across diverse sensor types.
-We hope this work opens up new possibilities for unifying wide-FoV segmentation.
+Our deformable decoder can seamlessly integrate with CNN-, transformer-, and Mamba-based backbones, imparting distortion-awareness capability to the overall model.
+To evaluate the effectiveness of our method, we conduct extensive experiments across five datasets for wide-FoV segmentation, including two 180° (WoodScape [62], SynWoodScape [45]) datasets and three 360° ( Stanford2D3D [1], Matterport3D [3], SynPASS [65]).
+Our framework performs consistently well across various datasets.
+Notably, while reducing the decoder parameters by $72\%$ and FLOPs by $97\%$, Deformable Mamba achieved state-of-the-art performance, with a ${+}2.5\%$ absolute improvement over previous VMamba [34] on the 360° Stanford2D3D dataset.
 To summarize, our contributions are as follows.
-  • We propose a Mamba-based framework termed Deformable Mamba, with the aim at unifying distortion-aware wide-FoV segmentation.
-  • A deformable decoder constructed with Deformable Mamba Fusion (DMF) module is specifically designed to boost the performance of mamba-based models in the wide-FoV segmentation task.
+  • We propose an efficient decoder based on the Mamba architecture, which addresses the gap in decoder-oriented Mamba-based efficient approaches.
+  • For the first time, we incorporate distortion-awareness exclusively at the Mamba decoder level, decoupling it from the overall framework, which enables integration with CNN-, Transformer-, and Mamba-based backbones to enhance distortion perception.
   • Extensive experiments on five datasets across a variety of indoor and outdoor, synthetic and real-world wide-FoV scenes, demonstrate the state-of-the-art performance of our method in wide-FoV segmentation.
+Figure 3: Overview of the Deformable Mamba (DMamba) framework.
+Given wide-FoV images (180° or 360°), the features extracted by an encoder, are fused by the proposed Deformable Mamba Decoder, which constructed by four Deformable Mamba Fusion (DMF) modules.
 2 Related Work
 2.1 Wide Field of View Segmentation
-Wide Field-of-View (Wide-FoV) semantic segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics.
-However, directly applying pinhole models [39, 2, 5, 15, 26, 65, 65, 36, 34, 72, 50, 58, 9, 19, 6, 70] to wide-FoV images often results in downgraded performance due to object deformations and image distortions.
+Wide-FoV image segmentation has received increasing attention as it enables broader scene understanding in applications such as autonomous driving and robotics.
+Wide-FoV images, like panoramic and fisheye images, offer expansive visual coverage but introduce unique challenges due to geometric distortions, spatial resolution variations, and uneven feature distribution.
+Early segmentation methods [36, 2, 5, 16, 25, 63, 63, 33, 32, 69, 47, 55, 10, 18, 6, 67] have achieved significant success on narrow-FoV image segmentation.
+However, directly applying pinhole models to wide-FoV images often results in downgraded performance due to object deformations and image distortions.
 Previous work [59, 60] propose the Panoramic Annular Semantic Segmentation (PASS) framework using a panoramic annular lens system.
@@
 Departing from these approaches, we are the first to propose a unified model applicable to all wide-FoV inputs, including panoramic and fisheye images.
-2.2 State Space Model
-CNNs [30] have achieved remarkable success in computer vision through effective local feature extraction and hierarchical representation learning capabilities, while Transformers [54] excel at modeling long-range dependencies despite their quadratic computational complexity.
-Recently, the state-space-model (SSM) based Mamba [16] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, ushering in a new paradigm for long sequence modeling.
-The emergence of Mamba in sequence modeling has naturally extended to vision tasks, Zhu et al.
-[75] adopt the patch embedding from Vision Transformer (ViT) [12] to project 2D image patches into sequential tokens compatible with Mamba architecture.
-To address the inherent discrepancy between image and sequence representations, a quadri-directional scanning and merging mechanism termed SS2D is introduced in [37] to supersede Mamba’s canonical 1D scan operation.
+2.2 Vision Mambas
+CNN [28] have achieved notable success in computer vision with efficient local feature extraction and hierarchical learning, while Transformer [51] excel at modeling long-range dependencies despite their quadratic complexity.
+Recently, the state-space-model (SSM)-based Mamba [17] architecture has emerged as a promising alternative, offering linear complexity and parallel computation through its selective scan mechanism, enabling efficient long-sequence modeling.
+The application of Mamba in sequence modeling has naturally extended to vision tasks.
+The serialized processing of input tokens in the Mamba architecture presents notable challenges when transferring to vision tasks.
+Zhu et al.
+[72] serialize the image using patch embedding [13], then scan the sequence from two directions.
+In [34], a quadri-directional scanning mechanism is proposed to supersede Mamba’s initial 1D scan operation to 2D scan.
 Subsequent work [24, 41, 23, 30, 66, 61, 4] introduce various scanning methods to explore variations of Mamba in visual representations.
-However, the study of using SSM for wide-FoV imagery is under-explored.
-Therefore, for the first time, we propose deformable Mamba to address both panoramic and fisheye semantic segmentation.
+However, these works have primarily leveraged the efficiency of the Mamba architecture from an encoder perspective, while its application in the decoder remains relatively constrained.
+Besides, while benefiting from the efficiency of the Mamba architecture, its application in wide-FoV dense prediction tasks remains unexplored.
+Therefore, for the first time, we propose a mamba-based decoder to address wide-FoV image segmentation.
 3 Methodology
-3.1 Preliminary
-Image Distortion.
-As shown in the left side of Fig.
-2, the point $A$ in the world coordinates undergoes extrinsic transformation to camera coordinates by Eq.
-(1), where $R$ and $T$ denote the rotation matrix and translation vector, respectively.
-Following Eq.
-(2), points in the camera coordinates are projected onto the image plane, with $f_{x}$ and $f_{y}$ representing the focal lengths and $c_{x},c_{y}$ defining the principal point.
-While straight lines are preserved in pinhole-approximate narrow-FoV cameras, only points within the sensor’s FoV (red box in Fig.
-2) are imaged, excluding points like $B^{\prime}$ that fall outside this region.
-Narrow-FoV cameras follow perspective projection principles where object size in the image varies inversely with distance between object and camera, introducing moderate distortion.
-$\begin{bmatrix}\mathbf{X}_{c}\\ \mathbf{Y}_{c}\\ \mathbf{Z}_{c}\end{bmatrix}=\mathbf{R}\begin{bmatrix}\mathbf{X}_{w}\\ \mathbf{Y}_{w}\\ \mathbf{Z}_{w}\end{bmatrix}+\mathbf{T},$ (1)
-$\begin{bmatrix}\mathbf{x^{\prime}}\\ \mathbf{y^{\prime}}\\ 1\end{bmatrix}=\begin{bmatrix}\mathbf{f}_{x}&0&\mathbf{c}_{x}\\ 0&\mathbf{f}_{y}&\mathbf{c}_{y}\\ 0&0&1\end{bmatrix}\begin{bmatrix}\frac{\mathbf{X}_{c}}{\mathbf{Z}_{c}}\\ \frac{\mathbf{Y}_{c}}{\mathbf{Z}_{c}}\\ 1\end{bmatrix}.$ (2)
-In contrast, wide-FoV cameras like fisheye cameras employ complex projection models involving light refraction.
-Kannala et al.
-[28] propose a generalized projection model for wide-FoV cameras, the distance between the image point and the principal point $r$ is approximated by:
-$\mathbf{r}(\mathbf{\theta})=\mathbf{k_{1}}\mathbf{\theta}+\mathbf{k_{2}}\mathbf{\theta}^{3}+\mathbf{k_{3}}\mathbf{\theta}^{5}+\mathbf{k_{4}}\mathbf{\theta}^{7}+\mathbf{k_{5}}\mathbf{\theta}^{9},$ (3)
-where $\theta$ is the angle between the principal axis and the incoming ray, $k_{i},i\in\{1,2,3,4,5\}$ are the parameters of wide-FoV camera model.
-While these cameras capture wider FoV, they suffer from more severe geometric distortions compared to the perspective projection of narrow-FoV cameras.
-As shown in the right side of Fig.
-2, barrel distortion in wide-FoV cameras causes straight lines to appear curved outward, particularly pronounced near image edges while the center remains relatively undistorted.
-This distortion manifests as radially-increasing displacement from the center, resulting in progressive spatial compression toward the image periphery.
-State Space Model. State space model (SSM), which describes the dynamics of a system, can be formulated as:
-$\begin{split}\mathbf{h}^{\prime}(\mathbf{t})&=\mathbf{A}\mathbf{h}(\mathbf{t})+\mathbf{B}\mathbf{x}(\mathbf{t}),\\ \mathbf{y}(\mathbf{t})&=\mathbf{C}\mathbf{h}(\mathbf{t}),\\ \end{split}$ (4)
-where $x(t)$ and $y(t)$ denote the inputs and outputs of the system, respectively, $h(t)$ is the inner states of the system, $A,B,C$ are the system parameters.
-Gu et al.
-[18] propose Structured State Space Sequence Model (S4) by initializing the parameters $A$ using Hippo Matrix (5) and discretizing parameters A and B by step size $\Delta$.
-$\mathbf{A}_{nk}=-\begin{cases}(2\mathbf{n}+1)^{1/2}(2\mathbf{k}+1)^{1/2}&\text{if }\mathbf{n}>\mathbf{k}\\ \mathbf{n}+1,&\text{if }\mathbf{n}=\mathbf{k}\\ 0,&\text{if }\mathbf{n}<\mathbf{k}.\end{cases}$ (5)
-In Mamba [16], a selective scan mechanism is proposed to further extend the SSM from time-invariant to time-variant, while also enabling parallel training.
-Observations.
-Based on HiPPO theory [17], Mamba [16] is inherently a sequence model, where the output of the current token only depends on preceding tokens, without considering subsequent ones - a limitation that becomes problematic for 2D images with inherent spatial structures.
-Unlike Transformer’s [54] mechanism in ViT [12], where attention computation for each patch considers all patches globally, Mamba processes image patches sequentially.
-To address this limitation, previous work [37, 75, 4, 76, 63] have explored different scanning patterns, including bidirectional or quadri-directional scanning approaches, partially addressing the challenges posed by Mamba’s unidirectional scanning nature in visual tasks.
-However, these multi-directional scanning approaches face another challenge.
-As illustrated in Fig 3., consider an input image embedded into 9 patches (labeled 1-9).
-When modeling patch 5 using SS2D’s [37] quadri-directional scanning mechanism, the scanning sequences are:
-  • (1) Left-Right: 1 - 2 - 3 - 4 - 5
-  • (2) Right-Left: 9 - 8 - 7 - 6 - 5
-  • (3) Top-Down: 1 - 4 - 7 - 2 -5
-  • (4) Down-Top: 9 - 6 - 3 - 8 - 5
-Figure 3: Visualization of SS2D quadri-directional scanning.
-Patch 1 and patch 3 maintain equal distances from patch 5, demonstrating uniform spatial distribution.
-As illustrated in sequences (1) (3) (4), the scanning trajectory from patch 1 to patch 5 traverses double the path length compared to the direct route between patch 3 and patch 5, despite their equivalent spatial distances.
-Since the Mamba model learns relationships between serialized tokens, the contribution of earlier patches in the scanning path to the current patch decreases as their distance from the current patch increases.
-Thus, along this scanning path, we infer that patch 1 is more important to patch 5 than patch 3.
-This priority is introduced by the scanning mechanism, but it is not a reasonable assumption.
-Additionally, when computing for other target patches, this priority changes based on the target’s position.
-Figure 4: Overview of the Deformable Mamba (DMamba) framework.
-Given wide-FoV images (180° or 360°), the features extracted by an encoder, are fused by the proposed Deformable Mamba Decoder, which constructed by four Deformable Mamba Fusion (DMF) modules.
 3.1 Overall Architecture
 An overview of our Deformable Mamba architecture is presented in Fig.
-4, which illustrates how we make Mamba-based models deformable and available for wide-FoV semantic segmentation.
+3, which illustrates how we make Mamba-based models deformable and available for wide-FoV semantic segmentation.
 It mainly follows the Encoder-Decoder architecture.
@@
 In our framework, there are two model sizes: DMamba-M (Mini) and DMamba-T (Tiny).
-More specifically, given the wide-FoV input image in the shape of $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$.
-The features extracted by the encoder undergo a hierarchical fusion process in the decoder.
+Specifically, given the wide-FoV input image of shape $H\times W\times 3$, a stem module is firstly utilized to patchify the original image into the shape of $\frac{H}{4}\times\frac{W}{4}\times 3$, the encoder gradually down-samples feature maps in the $l^{th}$ stage with channel dimensions $C_{l}\in\{96,192,384,768\}$ and resolutions $R_{l}\in\{\frac{H}{4}\times\frac{W}{4},\frac{H}{8}\times\frac{W}{8},\frac{H}{16}\times\frac{W}{16},\frac{H}{32}\times\frac{H}{32}\}$.
+These features are hierarchically fused in our proposed Deformable Mamba Decoder.
+The output features from the decoder’s final layer are first processed through a fully connected layer and subsequently transformed via feature channel convolutions to generate the predicted semantic map.
 3.2 Deformable Mamba Decoder
-The Deformable Mamba Decoder comprises four fusion and upsampling modules, where we introduce a novel Deformable Mamba Fusion (DMF) module for multi-scale feature integration.
-The initial DMF block processes dual inputs of ${\frac{H}{32}\times\frac{H}{32}\times 768}$ feature maps derived from the final encoder layer.
-The fused features are then processed through an upsampling module that simultaneously reduces channel dimensionality and increases spatial resolution, followed by channel mixing via an MLP layer.
-Following [49, 68, 35, 60], we utilize Pixel Shuffle [49] as the upsampling operation.
-In the subsequent stages, the second through fourth DMF blocks implement skip connections, facilitating interaction between the previously fused and upsampled features and their corresponding encoder features at matching scales.
-This progressive fusion strategy enables comprehensive multi-scale feature integration.
-The decoder’s output undergoes final processing through a fully connected layer and class mapping to generate the segmentation predictions.
-3.4 Deformable Mamba Fusion
-To address the aforementioned limitations of Mamba’s scanning mechanism in visual tasks (Sec.
-3.1), we propose the DMF module, which simultaneously adapts the features extracted by the Mamba-based encoder and integrates multi-scale features.
-As illustrated in the lower part of Fig.
-4, the DMF module takes two inputs: one is the feature $\mathbf{D}_{j}$ with dimensions $c\times h\times w$ output from the upper decoder layer, and the other is the corresponding encoder features $\mathbf{E}_{i}$ at matching scales, where $i$ and $j$ satisfy the relationship $i+j=5$.
-We employ the Cross-Scan [37] that scans $\mathbf{D}_{j}$ in four directions: (1) Left to Right, (2) Right to Left, (3) Top to Down, and (4) Down to Top, utilizing Mamba model’s global perception to generate four feature maps with resolution $c\times h\times w$ Following Cross-Merge [37], we first reverse features (2) and (4) along their last dimension and add them to (1) and (3) respectively, then transpose the $h$ and $w$ dimensions of feature (3) before adding it to (1).
-Inspired by previous works [10, 77, 56, 59], we propose a simple yet effective approach, utilizing deformable convolution [77] to enhance the features extracted by the Mamba model.
-Specifically, for a feature point in the encoder feature map $\mathbf{E}_{i}$, we attend to $\mathbf{K}$ neighboring feature points, allowing the model to learn offsets that compensate for the biases introduced by the scanning process.
-Additionally, convolution operations on these selected regions provide a more localized attention mechanism, further enhancing the visual representation capability of Mamba-based models in visual tasks.
-The mathematical formulation is as follows:
-$\mathbf{y}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{x}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (6)
-where $\mathbf{p}$ denotes the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively.
-And $\mathbf{m_{k}}$ is a learnable modulation scalar, which helps to modulate the input feature amplitudes from different spatial locations.
-For example, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
-The outputs from these two branches are concatenated and fed into a convolutional layer for feature fusion.
-This simple yet effective architecture successfully adapts the features extracted by the Mamba model, efficiently integrates multi-scale features, while endowing the entire decoder with deformable capabilities.
-Experiments on multiple datasets demonstrate both the effectiveness and efficiency of our proposed model.
+The Deformable Mamba Decoder comprises four efficient fusion and upsampling modules.
+Within this framework, we propose a novel Deformable Mamba Fusion (DMF) block, which effectively integrates multi-scale features while enabling the decoder to adaptively process distorted data.
+Considering the diverse types and degrees of distortion present in different wide-FoV data, our method is not tailored to any specific dataset.
+Instead, we incorporate distortion awareness into the model in a simple, broadly applicable manner.
+As shown in lower part of Fig.
+3, the DMF block takes two inputs, $E_{i}$ and $D_{j}$, where $E_{i}$ represents the multi-scale feature maps output by the encoder, and $D_{j}$ denotes the features fused by the DMF block.
+Here, indices $i$ and $j$ range from 1 to 4 and 0 to 3, respectively, following the constraint $i+j=4$.
+Specifically, for the first stage of DMF block, both inputs are derived from the feature maps of the last encoder layer, i.e., $E_{4}=D_{0}$.
+To maximize efficiency by minimizing decoder computational complexity, we employ only a single quadri-directional scanning layer, originally introduced in the 2D selective scan (SS2D) block by Liu et al.
+[34] for embedding feature $D_{j}$ into the state space.
+Figure 4: For an embedded 2D sequence, Mamba [17] employs a uni-directional scan from start to end (1).
+[72] utilizes a bi-directional scan combining (1) and (2), while [34] adopts a quadri-directional scanning incorporating (1), (2), (3), and (4).
+We posit that enabling model distortion-aware capability hinges on providing the model with sufficient deformable capacity, thereby endowing the model with the intrinsic ability to adaptively learn and generalize from distorted data representations.
+Deformable Convolutions (DCNs) [11, 74, 53, 56] render the rigid kernels in conventional convolution more flexible and dynamically adaptive by introducing data-dependent offsets, which allows the fixed sampling locations to be dynamically adjusted based on the data.
+[75] provides the Transformer-based model with a data-driven receptive field, enabling the model’s perception of image distortions and deformations.
+SSM, which describes the dynamics of a system, can be formulated as:
+$\begin{split}\mathbf{h}^{\prime}(\mathbf{t})&=\mathbf{A}\mathbf{h}(\mathbf{t})+\mathbf{B}\mathbf{x}(\mathbf{t}),\\ \mathbf{y}(\mathbf{t})&=\mathbf{C}\mathbf{h}(\mathbf{t}),\\ \end{split}$ (1)
+where $\mathbf{x}(\mathbf{t})$ and $\mathbf{y}(\mathbf{t})$ denote the inputs and outputs of the system, respectively, $\mathbf{h}(\mathbf{t})$ is the inner states of the system, $\mathbf{A},\mathbf{B},\mathbf{C}$ are the system parameters.
+Mamba [17] and its visual variants [72, 34, 24, 41, 23, 30, 66, 61, 4] are based on SSM.
+As shown in Fig.
+4, while methods such as [72, 4] have attempted to model the spatial information of images by expanding scanning paths, the inherent one-by-one token processing makes it more challenging to integrate a learnable deformation mechanism into state-space modeling, compared to CNN- and Transformer-based architectures.
+To address this, we adopt a compromise solution by following [65, 64, 68], we incorporate deformable convolution to augment the model’s deformable capacity.
+In particular, we feed the encoder feature $E_{i}$ into one single DCNv2 [74] layer with a kernel size of $\mathbf{3\times 3}$, which can be mathematically formulated as follows:
+$\mathbf{E_{out}}(\mathbf{p})=\sum^{K}_{k=1}\mathbf{w}_{k}\mathbf{m}_{k}\mathbf{E_{in}}(\mathbf{p}+\mathbf{p}_{k}+\Delta\mathbf{p}_{k}),$ (2)
+where $\mathbf{p}$ represents the current location in feature map, $\mathbf{w_{k}}$ and $\mathbf{p_{k}}$ denote the weight and initial offsets for the $\mathbf{k}$-th location, respectively.
+The parameter $\mathbf{m_{k}}$ is a learnable modulation scalar that modulates the input feature amplitudes from different spatial locations.
+In our implementation, $\mathbf{K}=9$ and $\mathbf{p_{k}}\in{(-1,-1),...,(1,1)}$ defines a $\mathbf{3\times 3}$ convolutional kernel with dilation 1.
+The outputs from both branches, each with dimensions $\{h\times w\times c\}$, are concatenated to form a tensor with dimensions $\{h\times w\times 2c\}$, which is subsequently processed by an upsampling module.
+To fully preserve the feature information from both branches while endowing the model with additional learnable capacity, we refrain from employing parameter-free upsampling methods such as bilinear interporlation.
+Instead, after an initial feature fusion via a series of $\mathbf{3\times 3}$ convolutions, we enhance the resolution of feature maps using the Pixelshuffle [46] operation, which can be formulated as:
+$\mathcal{PS}(T)_{h,w,2c}=T_{2h,2w,\frac{c}{2}},$ (3)
+where $\mathcal{PS}$ is an periodic shuffling operator that rearranges the elements of a input tensor $T$ from shape $\{h\times w\times 2c\}$ to $\{2h\times 2w\times\frac{c}{2}\}$.
+The upsampled features are linearly projected as the output of the DMF block.
+Uniquely, the feature map in the final stage, after undergoing convolutional fusion, is not subjected to the PixelShuffle operation for channel reduction and resolution enhancement.
+It is worth noting that in all the experiments presented in Sec.
+4, the decoder architecture is kept consistent, with no dataset-specific modifications.
+The only adjustment made is to the final output layer, which is tailored to accommodate the varying semantic categories across the datasets.
+Our Deformable Mamba Decoder leverages the advantages of the Mamba architecture’s linear computational complexity, resulting in superior computational efficiency.
+Simultaneously, it addresses the challenge that Mamba, in contrast to CNNs and Transformers, lacks the architectural ability to provide deformation awareness at the feature level.
+Besides, we introduce distortion awareness exclusively from the decoder perspective, decoupling it from the overall framework and thus enhancing the model’s flexibility and adaptability.
+Comprehensive qualitative and quantitative analyses in Sec.
+4 showcase the exceptional performance and efficiency of our decoder.
 4 Experiment Results
+To validate our deformable Mamba method, we compare with state-of-the-art approaches on three 360° panoramic and two 180° fisheye datasets.
+Additionally, we conduct experiments using CNN-, Transformer-, and Mamba-based architectures (e.g., ResNet, Swin, VMamba) to validate the efficiency and effectiveness of our Deformable Mamba Decoder.
 4.1 Datasets
-To study the unification of wide-FoV segmentation, we conduct experiments on three 360° and two 180° datasets.
 Stanford2D3D [1] consists of 1,413 indoor 360° images with a resolution of $512{\times}1024$, encompassing 13 distinct object categories.
 Matterport3D [3] has 10,615 indoor 360° images. Following [50, 21], 7,829 of panoramas with a resolution of 1024×2048 are used for training and 772 for evaluation.
-SynPASS [67] is a virtual outdoor 360° dataset from CARLA [13].
+SynPASS [65] is a synthetic outdoor 360° dataset from CARLA [14].
 It contains 9,080 images with a resolution of $1024{\times}2048$, covering various weather conditions (cloudy, foggy, rainy, sunny) and illumination (daytime, nighttime).
@@
 The learning rate schedule consisted of a linear warm-up for the first 1.5K iterations, followed by a polynomial decay with a power of 0.9.
+4.3 Effectiveness and Robustness analysis
+Table 1: Results on Stanford2D3D, using ResNet-50 [22] as the backbone with 23.5M parameters, 22.9G FLOPs and hierarchical embedding dimensions of {256, 512, 1024, 2048}.
+[…]
+mIoU Acc
+UperHead [54] (CVPR20) 40.5 250.7 53.34 62.42
+MusterHead [57] (CVPR21) 203.1 211.5 54.39 63.27
+CGRHead [38] (CVPR23) 282.6 31.4 54.12 62.94
+Ours 77.3 38.8 57.22 66.32
+Table 2: Results on Stanford2D3D, using Swin-T [35] as the backbone with 27.5M parameters, 25.6G FLOPs and hierarchical embedding dimensions of {96, 192, 384, 768}.
+[…]
+mIoU Acc
+UperHead [54] (CVPR20) 31.5 206.9 53.82 63.06
+MusterHead [57] (CVPR21) 19.1 21.4 52.36 54.54
+CGRHead [38] (CVPR23) 40.6 5.0 54.26 62.90
+Ours 11.2 6.0 55.09 65.30
+Table 3: Results on Stanford2D3D, using VMamba-T [34] as the backbone with 29.5M parameters, 25.3G FLOPs and hierarchical embedding dimensions of {96, 192, 384, 768}.
+[…]
+mIoU Acc
+UperHead [54] (CVPR20) 31.5 206.9 56.80 65.01
+MusterHead [57] (CVPR21) 19.1 21.4 58.12 66.97
+CGRHead [38] (CVPR23) 40.6 5.0 56.45 65.07
+Ours 11.2 6.0 59.30 68.99
+Adaptability to different backbones.
+As mentioned in Sec.
+1 and Sec.
+3, decoupling the deformable setting from the overall framework and introducing it exclusively from the decoder perspective somewhat reduces the model’s capability to perceive distortions in wide-FoV data.
+The advantage, however, lies in enhancing the model’s flexibility.
+Through the plug-in distortion-aware decoder, we enable the model to adaptively perceive image deformations for different backbones.
+On the Stanford2D3D panoramic dataset, we compared our Deformable Mamaba Decoder with UperHead [54], MusterHead [57], CGRHead [38], evaluating their adaptability to CNN-, Transformer-, and Mamba-based backbones.
+As shown in Tables 1, 2, and 3, compared to decoders lacking distortion perception, our method achieves the best performance across ResNet-50 [22], Swin-T [35], and VMamba-T [34] backbones, which demonstrates the significance of distortion perception for wide-FoV image understanding.
+Efficiency of the decoder.
+We emphasize that an efficient decoder is crucial for alleviating the overall model complexity.
+For instance, Liu et al.
+ultilize UperHead as the decoder to evaluate VMamba’s application in semantic segmentation tasks.
+As shown in Table 3, with VMamba as the backbone, the model has only 25.3G FLOPs, while the decoder UperHead, with 206.9G FLOPs, becomes the bottleneck of the models’ computational complexity.
+While MusterHead and CGRHead reduce computational complexity, they exhibit strong sensitivity to the model scale.
+As evidenced in Table 1, their parameter count experiences substantial growth with increasing embedding dimensions.
+In contrast, our proposed decoder showcases superior compatibility with diverse model scales while retaining lower temporal and spatial complexity.
+Based on the results, our method provides substantial distortion-aware capability for CNN-, Transformer-, and Mamba-based models while maintaining excellent efficiency, demonstrating its adaptability and robustness.
+4.4 Qualitative Results
+Visualization of distortion-aware capability.
+To showcase the distortion-aware capability of our decoder, we present a visual comparison against UperHead [54] lacking deformable desin on the 360° Stanford2D3D and 180° SynWoodScape datasets.
+As shown in Fig.
+5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings.
+In contrast, our proposed DMamba with Deformable Mamba Decoder effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
+Figure 5: Visualization of the wide-FoV segmentation results. From left to right: one 360° and two 180° segmentation.
+Analysis of feature maps.
+As shown in Figure 6, to illustrate how the DMF module enhances distortion awareness, we visualize the channel map extracted from the encoder and the corresponding output features of DMF block.
+Specifically, we present the results of channels #$\#6$ and #$\#68$ in the second layer of the first decoder stage.
+Prior to passing through the DMF block, the features primarily focus on global structures while lacking adequate emphasis on crucial regions such as small objects in cluttered indoor environments 6(a) or severely distorted areas in wide-FoV outdoor scenes 6(b).
+Without explicit distortion-aware adaptation, the feature representation remains less effective in preserving fine-grained spatial details and compensating for geometric distortions.
+The model shows enhanced attention to objects at different scales after traversing the DMF block.
+(a) Visualization of feature maps in indoor scenarios
+(b) Visualization of feature maps in outdoor scenarios
+Figure 6: Visualization of feature maps.
+From left to right: features extracted by Mamba-based encoder (Before DMF) and features fused by the DMF module (After DMF).
+Effective detection and recognition of small objects in indoor scenarios (a), such as pictures on walls, requires the model to exhibit enhanced local perception capabilities.
+In contrast, outdoor scenarios (b) captured by wide-FoV cameras are often characterized by significant object deformation and distortion.
+To facilitate accurate scene understanding in such cases, the model must be equipped with a robust ability to handle and perceive distortion effectively.
 4.5 Quantitative Results
-Given that previous works were largely constrained to specific Field-of-View (FoV) settings, which poses challenges for fair cross-dataset evaluation.
-Therefore, we selected representative state-of-the-art models as our baselines, including SegNeXt [21], which achieves superior performance on pinhole images, and 360SFUDA [73], which is specifically designed for panoramic scenarios.
-To ensure fair comparison, we carefully match the parameter counts across all implementations and conduct extensive evaluations on the aforementioned datasets with varying FoV characteristics.
+We extensively evaluate our proposed method across the five aforementioned datasets, addressing the critical challenge of data heterogeneity prevalent in existing methods.
+Through meticulous comparative experiments with state-of-the-art approaches and a careful re-implementation of [20, 70], we systematically investigate the performance variations induced by diverse data distributions and types, thereby providing comprehensive insights into model generalizability and robustness.
 Table 4: Results on 360° segmentation datasets of Stanford2D3D (S2D3D) fold-1 and 360FV-Matterport (MP3D). denotes our re-implementation.
@@
 Trans4PASS+ [65] (PAMI24) Trans4PASS+ 54.0 42.6
-SegFormer-B2 [58] (NeurIPS21) MiT-B2 51.9 45.5
+SegFormer [55] (NeurIPS21) MiT-B2 51.9 45.5
 360BEV [50] (WACV24) MiT-B2 54.3 46.3
@@
 DMamba-T (Ours) VMamba-S 60.2 49.3
-Results on 360° Stanford2D3D.
-Table 1 shows superior performance of our method compared to both panoramic [14, 51, 74, 66, 67, 53, 33, 73] and pinhole-based [21, 37] methods.
-Specifically, our model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines.
-Notably, compared to the VMamba-T + UperNet [57], our lightweighted DMamba-M with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing parameter count by 30% and FLOPs by 80%.
-Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba-T + UperNet, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
+Results on 360° Stanford2D3D dataset.
+Table 4 shows superior performance of our method compared to both distortion-aware [15, 48, 71, 64, 65, 50, 31, 70] and non-distortion-aware [20, 34] methods.
+Specifically, our DMamba-M model with the VMamba-T as backbone achieves a significant mIoU improvement of ${+}2.9\%$ to ${+}13.7\%$ over panoramic baselines.
+Notably, compared to the VMamba [54], our lightweighted decoder with the same backbone achieves a ${+}2.5\%$ higher mIoU while reducing overall parameters by 20% and FLOPs by 86%.
+Furthermore, our DMamba-T model, which has a comparable parameter count to VMamba, demonstrates even better performance with a ${+}3.4\%$ mIoU improvement.
+Results on 360° Matterport3D dataset.
+While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of ${+}0.3\%$ mIoU over the previous SOTA model SegNeXt as shown in Table 4.
+Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
 Table 5: Results on 360° segmentation dataset SynPASS. denotes our re-implementation.
@@
 DMamba-T(Ours) 42.78 84.24 22.57 00.25 49.20 33.25 76.85 98.56 62.05 68.79 82.91 02.11 32.35 98.52 02.42 01.73 30.68 50.46 37.40 32.04 05.85 21.59 46.32
-Results on 360° Matterport3D.
-While our DMamba-M demonstrates remarkable performance on Stanford2D3D dataset, the challenge becomes more pronounced on large-scale Matterport3D panoramic datasets, where we observe a relatively modest improvement margin of 0.3 mIoU over the previous SOTA model SegNeXt as shown in Table 1.
-Remarkably, our DMamba-T achieves more substantial gains of ${+}0.9\%$ and ${+}1.7\%$ mIoU compared to SegNeXt and VMamba-T respectively, with comparable model capacities.
-Results on 360° SynPASS.
-Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the SynPASS dataset.
-It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset.
-This discrepancy underscores the sensitivity of certain architectures to distribution shifts.
-In contrast, our method, as shown in Table 2, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods.
-Specifically, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic.
-For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
 Table 6: Results on 180° segmentation dataset WoodScape. denotes our re-implementation.
@@
 DMamba-T (Ours) 83.90
-Results on 180° WoodScape.
-In Table 3, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [64].
-Semi-supervised methods such as MeanTeacher [52], CPS [7], and FishSegSSL [43] show lower performance due to the limited data.
-Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [31] model.
-Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively.
-The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
 Table 7: Results on 180° segmentation dataset SynWoodScape. denotes our re-implementation.
@@
            DMamba-T (Ours)            69.62
-Results on 180° SynWoodScape.
+Results on 360° SynPASS dataset.
+Beyond real-world datasets, we evaluate the generalizability of our method on synthetic and outdoor scenes on the 360° SynPASS dataset.
+It is worth noting that despite SegNeXt’s superior performance over 360SFUDA in real-world scenarios, it exhibits a substantial performance degradation on the synthetic dataset.
+This discrepancy underscores the sensitivity of certain architectures to data distribution shifts.
+In contrast, our method, as shown in Table 5, achieves the state-of-the-art mIoU of 42.78%, surpassing all existing panoramic- and pinhole-based methods.
+Explicitly, our DMamba-T overall achieves top scores on pedestrian, pole, traffic sign, static and dynamic.
+For traffic sign with slender properties, our model enjoys more than ${+}6.0\%$ mIoU.
+Results on 180° WoodScape dataset.
+In Table 6, we present a comparison of our method with other models using 180° fisheye inputs from the WoodScape dataset [62].
+Semi-supervised methods such as MeanTeacher [49], CPS [8], and FishSegSSL [40] show lower performance due to the limited data.
+Compared to multi-task models, our proposed DMamba, designed with deformable settings, outperforms all variants of the OmniDet [29] model.
+Although SegNeXt and 360SFUDA perform well on the WoodScape dataset, our method, with a more compact architecture still achieves a higher mIoU by ${+}0.9\%$ and ${+}0.8\%$, respectively.
+The DMamba-T variant further advances the state-of-the-art mIoU to 83.9%, confirming DMamba’s effectiveness in capturing heavily distorted object information in wide-FoV segmentation.
+Results on 180° SynWoodScape dataset.
 Due to the lack of baseline comparisons on SynWoodScape, we limited our evaluation to SegNeXt and 360SFUDA.
-Similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic 180° data.
-This further confirms the challenge of achieving consistently high performance across diverse datasets.
+As shown in Table 7, similar to the results on SynPASS, SegNeXt and 360SFUDA perform similarly on real 180° datasets, while the former exhibits a significant performance degradation on synthetic data.
+This further confirms the challenge of achieving consistently high performance when data distribution shifts.
 In comparison, our DMamba-M of similar scale outperforms them by ${+}4.36\%$ and ${+}1.73\%$ mIoU, respectively, achieving an improved segmentation accuracy of 69.20 mIoU on SynWoodScape.
 These consistent improvements across all five wide-FoV datasets demonstrate the effectiveness of our proposed framework in unifying wide-FoV segmentation.
-Figure 5: Visualization of the wide-FoV segmentation results. From left to right: one 360° and two 180° segmentation.
-4.4 Qualitative Results
-To demonstrate the improved performance of our model over Mamba-based models, we compared the visualization results of DMamba and VMamba on the Stanford2D3D and SynWoodScape.
-As discussed in Sec.
-3.1 and  3.4, the scanning approach in Mamba-based models introduces biases in visual tasks that limit their effectiveness in dense prediction tasks.
-These biases increase randomness and reduce the model’s ability to capture local spatial relationships, which is particularly evident in the recognition and segmentation of smaller objects.
-As shown in Fig.
-5, the baseline model struggles to accurately segment narrow and elongated areas such as door frames, vegetation contours, traffic light poles, and lane markings.
-In contrast, our proposed DMamba effectively addresses these challenges, significantly improving segmentation performance on small and elongated objects.
-Table 5: Ablation study for analysis of backbones (BB) and decoders (Dec) on the 360° dataset Stanford2D3D.
-[…]
-BB
-Dec
-Total
-BB
-Dec
-Total
-mIoU
-mAcc
-ResNet-UperNet ResNet50 [23] UperHead [57] 44.3 23.5 40.5 64.0 22.9 250.7 273.6 53.34 62.42
-ResNet-UperFormer ResNet50 [23] MusterHead [60] 14.3 23.5 203.1 226.6 22.9 211.5 234.4 54.39 63.27
-ResNet-CGRSeg ResNet50 [23] CGRHead [41] 58.0 23.5 282.6 306.1 22.9 31.4 54.3 54.12 62.94
-ResNet-DMamba (Ours) ResNet50 [23] DMambaHead 27.9 23.5 77.3 100.8 22.9 38.8 61.7 57.22 66.32
-Swin-UperNet Swin-T [38] UperHead [57] 34.9 27.5 31.5 59.0 25.6 211.5 237.1 53.82 63.06
-Swin-UperFormer Swin-T [38] MusterHead [60] 30.2 27.5 19.1 46.6 25.6 21.9 47.5 52.36 54.54
-Swin-CGRSeg Swin-T [38] CGRHead [41] 52.2 27.5 40.6 68.1 25.6 5.5 31.1 54.26 62.90
-Swin-DMamba (Ours) Swin-T [38] DMambaHead 37.3 27.5 11.2 38.7 25.6 6.5 32.1 55.09 65.30
-VMamba-UperNet VMamba-T [37] UperHead [57] 35.0 29.5 31.5 61.0 25.3 206.9 232.2 56.80 65.01
-VMamba-UperFormer VMamba-T [37] MusterHead [60] 29.8 29.5 19.1 48.6 25.3 21.4 46.7 58.12 66.97
-VMamba-CGRSeg VMamba-T [37] CGRHead [41] 42.2 29.5 40.5 70.0 25.3 5.0 30.3 56.45 65.07
-DMamba-M (Ours) VMamba-T [37] DMambaHead 40.0 29.5 11.1 40.6 25.3 6.0 31.3 59.30 68.99
-DMamba-T (Ours) VMamba-S [37] DMambaHead 17.9 49.4 11.1 60.5 44.7 6.0 50.7 60.18 71.68
 4.6 Ablation Study
-Effect of Backbone.
-In Table 5, we ablate three different kinds of commonly-used model backbones, namely CNN-based ResNet [23], ViT-based Swin Transformer [38] and Mamba-based VMamba [37].
-For the sake of a fair comparison, we choose the backbones with a comparable number of learnable parameters and FLOPs.
-According to Table 5 it can be observed that the models equipped with VMamba [37] outperform other models consistently while CNN-based ResNet50 and ViT-based Swin Transformer achieve similar performance.
-This conclusion can be validated by employing the same decoder, such as UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead, along with a consistent model backbone to perform the segmentation task.
-Notably, the combination of the VMamba [37] backbone with DMambaHead, resulting in our Deformable Mamba model, demonstrates superior performance compared to other variant combinations by a maximum of over $+6.9\%$ mIoU.
-Effect of Decoder.
-Apart from the backbones, we also ablate the decoders.
-For every backbone, we conduct four experiments with different decoders, i.e., UperHead [57], MusterHead [60], CGRHead [41], and our proposed DMambaHead.
-All results are reported in Table 5.
-Using the same model backbone, our proposed DMambaHead outperforms other widely-used decoders by a $+3\%$ mIoU while a complexity degradation of $+20\%$, making the Deformable Mamba and its variants extremely practical in real-time wide-FoV accurate segmentation.
-It’s worth noting that the DMambaHead with VMamba-T [37] backbone achieves a drop of FLOPs from $232.2$G to $31.3$G, a drop of the learnable parameters from $61$M to $40.6$M while the performance of Deformable Mamba increases from a $+56.8\%$ to $+59.3\%$ mIoU compared with the VMamba-T [37] and UperHead [57] combination.
-The reason behind this is that our proposed DMambaHead is specifically designed for wide-FoV segmentation.
-Our deformable design enables the model to focus on the distortion introduced by the wide-FoV fisheye or panoramic images.
-This ablation proves that not only the VMamba [37] backbone but also the proposed DMambaHead contributes to the superior performance and efficiency of the Deformable Mamba model.
-(a) Results on Matterport3D.
-(b) Results on Stanford2D3D.
-Figure 6: Analysis of segmentation with different wide FoVs.
-Analysis of FoVs.
-To further analyze the wide-FoV segmentation, we conduct omni-directional analysis.
-As shown in Fig.
-6, there are six different FoVs, ranging from 60° to 360°.
-Compared to the baseline VMamba [37], our DMamba methods obtain relative better results on all different FoVs.
-Note that the 60° performance of all methods are lower than other degrees on the Stanford2D3D dataset.
-One reason is that the object distribution and room layout in the 60° direction are different between Stanford2D3D and the Matterport3D datasets.
-Nonetheless, our Mini- and Tiny-version methods show consistent improvements over the baseline on both datasets in different FoVs.
+To further analyze the effect of different components on our proposed framework, we conduct ablation study on the Stanford2D3D dataset.
+Analysis of scanning mechanism.
+To ensure consistency, we adopt the same scanning method as in [34], as our model primarily uses its Mamba-based encoder.
+As mentioned in Sec.
+1, the SSM-based Mamba processes input tokens sequentially, which is suitable for textual data.
+However, for image data, where tokens have spatial relationships, multi-directional scanning is more appropriate.
+To highlight the difference between various scanning methods on our framework, we conducted ablation experiments on the Stanford2D3D dataset.
+As shown in Table 8, both uni-directional scanning and bi-directional scanning results in slight performance drop.
+Table 8: Analysis of scanning methods.
+        Scan Method       mIoU       mAcc
+      Uni-direction       57.53       67.52
+      Bi-direction       57.94       67.66
+      Quadri-direction (Ours)       59.30       68.99
+Study of defomable convolution.
+In Sec.
+1 and Sec.
+2, we outlined the limitations of tightly coupling deformable designs with the entire model and the challenges faced by Mamba-based models in integrating deformable mechanisms due to their sequential processing of input tokens.
+We propose to introduce sufficient distortion awareness into the model by paralleling DCN in the decoder and integrating it through feature fusion.
+We replace deformable convolution with traditional one as shown in Table 9, the results prove the effectiveness of this trade-off solution.
+Table 9: Ablation Study of deformable designs.
+  Deformable mIoU mAcc
+57.45 66.52
+59.30 ($+1.85\%$) 68.99 ($+2.47\%$)
+Analysis of upsampling methods.
+Sampling methods play a critical role in our approach.
+As mentioned earlier, our decoder utilizes two branches to achieve complete feature modeling and the incorporation of distortion awareness respectively.
+To optimize computational efficiency, we adopt a single-layer module in each branch, which reduces model complexity while limiting the model’s learnable capacity.
+As a result, we disregard parameter-free interpolation-based upsampling methods, opting instead to fuse features from both branches and employ periodic arrangement to preserve all information.
+In Table 10, we compare bilinear and bicubic methods, showing the importance of our design.
+Table 10: Analysis of upsample methods.
+        Upsample Method       mIoU       mAcc
+      Bilinear Interpolation       56.90       66.08
+      Bicubic Interpolation       56.97       66.12
+      PixelShuffle (Ours)       59.30       68.99
 5 Conclusion
@@
 Most existing models are tailored for pinhole camera images, making it challenging to extend their success to wide-FoV domains due to the severe image distortion and object deformation.
-Unlike previous approaches that require a specific architecture for each FoV or dataset, we propose a unified framework, Deformable Mamba, built on the emerging State Space Model (SSM) mechanism.
-Our framework is constructed with a flexible deformable decoder, which can be easily adapted to various models and datasets.
-Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our unified framework effectively addresses the unique challenges of wide-FoV segmentation.
+We incorporate the model’s distortion-aware capability exclusively from the decoder perspective, decoupling it from the overall architecture to enhance flexibility and facilitate seamless integration with various model architectures.
+Consistent improvements across five diverse wide-FoV datasets, ranging from indoor to outdoor, from synthetic to real-world scenes, demonstrate that our method effectively addresses the unique challenges of wide-FoV segmentation.
 Limitations and Future Work.
-Using large language models (LLMs) for unifying the wide-FoV segmentation is a potential solution but not discussed in this work, while LLMs can further include the segmentation methods that use vision-language architectures.
-Besides, cross-domain evaluation, e.g., transferring from a synthetic to a real wide-FoV scene, or from a narrow-FoV camera to a wide-FoV camera, will be further explored in our future work.
+Given the rapid advancement of Large Language Models(LLMs), integrating multi-modal LLMs presents a promising direction for future research.
+Their strong generalization capabilities and semantic understanding could significantly enhance our unified semantic segmentation framework, particularly in handling diverse wide-FoV segmentation tasks.
+This integration could bridge the gap between linguistic semantic understanding and geometric feature learning, potentially leading to more robust and versatile segmentation models.